### v1

# Midichlorian-assisted cold fusion Confidence–Diversity Calibration of AI Judgement Enables Reliable Qualitative Coding

Abstract
LLMs enable qualitative coding at large scale, but assessing the reliability of their output remains challenging in domains where human experts seldom agree. Analysing 5 680 coding decisions from eight state-of-the-art LLMs across ten thematic categories, we confirm that a model’s mean self-confidence already tracks inter-model agreement closely (Pearson ). Adding model diversity—quantified as the normalised Shannon entropy of the panel’s votes—turns this single cue into a dual signal that explains agreement almost completely (). The confidence–diversity duo enables a three-tier workflow that auto-accepts 35 % of segments with <5 % audit-detected error and routes the remainder for targeted human review, cutting manual effort by up to 65 %. Cross-domain replication on six public datasets spanning finance, medicine, law and multilingual tasks confirms these gains ( improvements of 0.20–0.78). Our results establish a generalisable, evidence-based criterion for calibrating AI judgement in qualitative research.
Introduction
Qualitative research has long faced a structural trade–off between scale and reliability. Traditional workflows typically depend on two–coder adjudication, yet resulting Cohen’s often remains only “moderate” while labour costs grow super–linearly with corpus size.1 Yet in domains lacking a single ground truth, researchers still need principled criteria for deciding when an AI label can be trusted. Drawing on calibration research in metacognitive monitoring, we treat self–confidence as a first–order cue and complement it with panel diversity, an error signal that reveals collective blind spots.2 From a human–behaviour perspective, our study therefore centres on how analysts allocate attention within human–AI collaboration, translating cognitive signals into workflow decisions that optimise expert effort. Computational text-analysis methods—ranging from early topic models to state-of-the-art large language models (LLMs)—are rapidly entering the methodological toolkit of social scientists and health researchers, promising to automate labour-intensive qualitative tasks such as open coding and theme extraction 3, 4, 5, 6, 7, 8. By partitioning the confidence–diversity space into actionable zones, we created a tiered review system that efficiently allocates human attention.
Indeed, empirical analyses and methodological reviews consistently show that inter-coder reliability in qualitative studies often falls only in the “moderate” range even among trained researchers1, 9. Disagreement reflects interpretive uncertainty, disciplinary lenses and contextual nuances that resist reduction to binary truth values10, 11. This makes it difficult to judge when an AI-generated label should be trusted, because standard accuracy metrics presuppose an authority against which the model can be scored.
Most recent work therefore turns to metacognitive signals. Self-reported confidence has emerged as a convenient proxy: if a model “knows that it knows”, high-confidence predictions might be prioritised while low-confidence ones are routed for review. However, modern neural networks are notorious for poorly calibrated probabilities12, 13, and confidence alone cannot expose blind spots shared by an otherwise unanimous ensemble of models.
Here we propose to complement confidence with model diversity—the dispersion of predictions across differently trained or prompted LLMs. Diversity has long served as an error bound in ensemble learning14, yet remains under-explored in qualitative research. We analyse eight state-of-the-art LLM variants—four single-response models and their chain-of-thought counterparts—that each coded 71 interview segments against a ten-category framework, yielding 5 680 decisions. A ninth model is introduced only in cross-domain validation (see Methods). Our approach focuses on relatively straightforward coding tasks with conceptually accessible categories that are interpretable to human coders, establishing a foundation for reliable quality assessment. We show that mean self-confidence explains much of the observed inter-model agreement (Pearson ), and that adding model diversity produces a dual indicator that predicts agreement almost perfectly ().
Finally, we show that this confidence–diversity principle generalises across six public datasets spanning finance, medicine and law, and that its predictive accuracy remains robust to the exact weighting of the two signals: although an in-sample grid search favours , the intuitive 0.6/0.4 split performs within 0.03 MAE across domains.
Related Work
At the intersection of qualitative content analysis, model–calibration theory and ensemble–uncertainty estimation, three strands of scholarship inform our study.
Reliability in qualitative coding. Classic content–analysis texts demonstrate that interpretive categories must be both conceptually coherent and empirically reliable. Formal checks usually rely on chance–corrected agreement metrics such as Cohen’s or Krippendorff’s ; empirical studies and methodological guidelines consistently classify values above as “almost perfect” agreement—and report that such cases are uncommon in practice.15, 16, 11, 17, 18 Recent work therefore explores whether computational models—from earlier topic-modelling pipelines to contemporary LLMs—can assist human coders at scale while preserving rigour, reporting promising early results in political science, communication and social research.6, 3, 4 Yet even with multiple annotators, majority voting improves stability only marginally because humans (and often models) share similar blind spots.
Calibration of modern neural networks. Large–scale studies reveal that soft–max probabilities produced by deep classifiers are often miscalibrated; simple post–hoc schemes such as temperature scaling shrink but do not eliminate over–confidence.12, 19, 13 Broader surveys propose dedicated metrics to quantify calibration quality.20 Deep ensembles add epistemic uncertainty, with Bayesian approximations such as SWAG further refining predictive distributions.21, 22 However, all these methods presuppose access to an oracle ground–truth label, a luxury rarely available in exploratory qualitative research.
Ensemble voting and uncertainty in NLP. Disagreement among independently prompted or fine–tuned models has been leveraged to flag hard examples, measure trust scores, and reduce hallucinations.23 Recent work on uncertainty–aware majority voting extends this logic to crowdsourced labels.24 Diversity within the ensemble therefore serves as an additional uncertainty cue that complements confidence; classic machine–learning theory offers quantitative diversity indices.25 Our study unifies confidence and diversity in the absence of ground truth, showing that their linear combination predicts agreement with , and applies this mechanism to qualitative coding for the first time.
Results
Human baseline comparison
To contextualise the dual–signal gains we first benchmarked model performance against a traditional two–coder baseline and our full three–tier review workflow (Table 1). Across the 5 680 coding decisions the eight–model majority already surpasses typical human reliability, achieving a chance–corrected Cohen’s and halving the error rate relative to double human coding. Integrating the dual–signal triage and expert reconciliation lifts effective reliability to while trimming the residual error rate to 4.1%. These figures derive from the audited interview corpus of 71 segments (see Methods).
| Condition | Cohen’s | Error rate (%) | Decisions () |
| Human pair (research assistants) | 0.67 | 11.4 | 1 420 |
| AI majority (8 models) | 0.89 | 6.3 | 5 680 |
| Three–tier workflow (AIexpert) | 0.93 | 4.1 | 5 680 |
Dual–signal mechanism eliminates residual disagreement
We first investigated whether an LLM’s mean self-confidence () is a reliable proxy for inter-model agreement (). Our analysis spans 5 680 coding decisions produced by eight state-of-the-art models on 71 interview segments. The eight-model panel achieves an accuracy of 93.7 %, a chance-corrected Cohen’s , and a strong correspondence between mean confidence and agreement (, ).
For completeness, we formalise the two quantities as
where denotes the number of affirmative votes among the eight models. Hence with indicating unanimity. Model diversity is computed as the normalised Shannon entropy of the vote distribution
which maps perfect consensus to and a balanced split (4–4) to .
These residuals correlate strongly with model diversity (), defined as the normalised entropy of the eight binary votes. Categories with greater diversity are under-predicted by the confidence-only model, whereas near-consensus categories are over-predicted (Fig. 1B; ). Combining and yields a dual-signal regression plane that captures almost all variance (; Fig. 1C), eliminating bias without additional features. While item‐level cross-validation identifies a pure‐diversity optimum (; Methods, Extended Data Fig. 7), we deliberately adopt the intuitive split throughout the paper because it transfers more reliably to six external datasets (median MAE) and aids substantive interpretation. Representative high-diversity examples are provided in Supplementary Table 5.
To systematically leverage this dual-signal pattern, we developed a mathematical model that formalizes the relationship between confidence, diversity, and agreement. This led us to an optimized regression equation that quantifies their interaction.
Enhanced regression unifies confidence and diversity
Building on the single-cue analysis, we fitted a multiple linear regression that combines average confidence and model diversity to predict agreement:
At the category level this two-variable model explains of the variance—an improvement of 12 percentage points over the confidence-only baseline (). The mean absolute error drops from 3.06 % to 1.48 %, confirming that diversity corrects systematic biases left by confidence alone. Full coefficient estimates with classical t–based 95 % confidence intervals are reported in Supplementary Table 10; percentile bootstrap intervals based on 1,000 resamples (Supplementary Table 11) show the same directional pattern, confirming that the diversity term remains significantly negative despite the small sample size. An item-level analysis over all 710 coding decisions (Supplementary Table 12) produces virtually identical slopes under both OLS and mixed–effects specifications, providing further evidence of model robustness across levels of aggregation.
Having established the predictive power of our dual-signal model, we next developed a practical workflow to translate these insights into actionable decision rules for human-AI collaboration.
Three-tier workflow reduces manual review
Leveraging the dual indicator, we mapped every category into a confidence–diversity plane and partitioned it into three actionable zones (Fig. 2A). We operationalised a quantitative risk score
This 0.6/0.4 weighting was chosen for interpretability and cross-domain robustness (see Methods). Despite the data-driven optimum lying at (see Methods), we retain the 0.6/0.4 split to balance interpretability and cross-domain stability (external ).
Among the ten thematic categories analysed here, one falls in the green zone, four in amber and five in red, translating to an estimated 35% of coding points that can be automatically accepted, 15% requiring only a lightweight audit, and 50% necessitating full adjudication. Crucially, this workflow shows that errors are highly concentrated: green-zone items exhibit residual error, whereas red-zone items concentrate 87 % of all disagreements (Fig. 2B). Residual error here denotes the proportion of auto-accepted items whose AI label was overturned during the 20 % expert audit ( green-zone samples). This asymmetric error distribution—where 87% of all errors concentrate in just 50% of the data (Extended Data Fig. 5)—is the key mechanism that enables substantial efficiency gains without compromising reliability.
Figure 2C illustrates the distribution of coding points across risk categories based on our quantitative risk score. The horizontal dashed line indicates the proportion of points selected for expert audit (4.7%), demonstrating how our approach significantly reduces the manual review burden while maintaining high reliability. The efficiency-accuracy trade-off curve (Figure 2D) shows how different confidence thresholds affect both the pass rate (blue line) and error rate (red dashed line). Our selected threshold of 4.5 (vertical dotted line) achieves an optimal balance between minimizing errors and maximizing automation.
Expert adjudication confirms workflow reliability
To validate the three-tier workflow against human expert judgment, we conducted a targeted expert audit on the most challenging segments. We identified 266 instances where AI models disagreed with original human codes and submitted these for independent review by three domain experts. This adversarial sampling approach deliberately focused on potential AI errors rather than examining random segments, providing a conservative test of AI reliability.
The expert panel achieved substantial inter-rater agreement (Fleiss’ 26 ) and produced a consensus judgment for each contested segment. In 87.6% of these challenging cases, the AI panel’s majority decision was confirmed as correct by expert consensus, indicating that the original human codes—not the AI judgments—contained errors (Extended Data Fig. 4). For a narrative walkthrough of one such correction, see Supplementary Note 1. This finding reduced the estimated gap between AI and optimal human coding by 67%, suggesting that well-calibrated LLMs can match or even exceed the reliability of typical human coders in qualitative analysis. Additional case studies illustrating ambiguity resolution and AI false positives are provided in Supplementary Notes 2 and 3, respectively.
Error analysis (Extended Data Fig. 5) revealed that remaining disagreements clustered around three categories: ambiguous segment boundaries (41%), genuine interpretive differences (37%), and conceptual edge cases (22%). These patterns inform our recommendations for human-AI collaborative coding (see Discussion), particularly the need for clear segment delineation and concept definitions prior to large-scale AI deployment.
Prompting style and model family performance
Figure 3 benchmarks four prominent LLM families—Claude, GPT, DeepSeek and Gemini—under two prompting styles: the default single–response (S) and an explicit chain–of–thought variant (T). Across the full 5 680–decision panel the two styles reach virtually identical agreement (single: 90.4 %, thinking: 90.6 %, paired ) and near–overlapping confidence distributions (mean vs. ) 27, 28.
Family–level patterns are more pronounced. Mean Cohen’s ranks Claude highest at 0.92, followed by GPT (0.90), DeepSeek (0.88) and Gemini (0.85). The ordering is consistent under both prompting styles, suggesting that architectural and training data differences, rather than the presence of chain–of–thought, dominate reliability. Confidence mirrors the same ladder (Fig. 3B), reinforcing that better–calibrated families also achieve stronger inter–model agreement.
Bootstrap resampling confirms Claude’s lead and shows no meaningful difference between single-response and chain-of-thought variants (Supplementary Table 4).
A comprehensive, multi‐panel breakdown of these approach‐level patterns—including category‐specific differences, Jaccard similarity of positive codes, and confidence–agreement calibration—is presented in Extended Data Fig. 3.
The performance hierarchy remained consistent across all ten thematic categories, with Claude maintaining its lead in nine categories and tying with GPT in one (Fig. 3). This consistency suggests that architectural differences and training data quality, rather than domain-specific optimizations, drive overall reliability in qualitative coding tasks. Notably, Claude exhibited the smallest gap between confidence and actual agreement (), indicating superior calibration—a critical factor when deploying LLMs in human-AI collaborative workflows.
Taken together, these results advise practitioners to prioritise high‐quality single‐response models and to reserve chain‐of‐thought prompting for tasks where transparency, not throughput, is paramount. In subsequent analyses we therefore rely on single‐response variants unless explicitly noted.
Cross-domain validation across six public datasets
With the three-tier workflow validated on our interview dataset, we next examined whether this approach would generalize to different domains and task types. We applied the dual indicator to six external datasets spanning diverse domains: financial sentiment analysis, medical diagnosis coding, legal case classification, moral reasoning assessment, machine translation quality evaluation, and multilingual natural language inference (600 sentences each). This cross-domain testing is critical to establish whether the confidence–diversity mechanism generalizes beyond our original qualitative interview context.
Figure 4 demonstrates that our findings robustly transfer across all tested domains. For five of the six corpora we used a four–model single–response panel (DeepSeek–S, GPT 4o–S, Gemini–S and Llama4–Maverick); only the biomedical PubMed–RCT benchmark retained the chain–of–thought variants, yielding an eight–model panel. Average confidence continues to track agreement closely (mean Pearson across domains), while adding diversity substantially elevates the explanatory power. Across five datasets the dual-signal model lifts from baseline values of – up to – (Supplementary Table 3), a jump exceeding 50 percentage points. The Finance–News set is a notable outlier: its binary sentiment task is already near ceiling, with 83 % of sentences receiving unanimous labels and average confidence . Coupled with the smaller four-voter panel that allows only two non–zero diversity levels, the enhanced regression can augment only marginally (0.146 0.154). The confidence–diversity calibration yields Cohen’s improvements ranging from (for machine translation) to (for legal classification), with a mean gain of across all datasets.
All improvements were statistically significant at using 5,000 bootstrap resamples, confirming that these substantial gains are not attributable to chance. The consistent performance across this diverse range of tasks—from subjective moral reasoning to technical legal classification—underscores the generality of the confidence–diversity principle for calibrating AI judgements in qualitative analysis.
Discussion
Summary of Key Findings
Our findings establish three headline results. (i) A simple dual indicator—mean confidence plus vote diversity—explains 98% of variance in inter-model agreement, doubling the accuracy of confidence alone. (ii) Routing coding points by this indicator enables a three-tier workflow that auto-accepts one-third of decisions and cuts human effort by 65% while preserving accuracy. (iii) Both expert adjudication and tests on six external datasets confirm that these gains generalise beyond the original interview corpus. The distribution of errors across risk tiers represents a particularly important finding: 87% of all errors concentrate in just 50% of the data (the high-risk tier), while the low-risk tier accounts for 35% of data but only 5% of errors. This asymmetric error concentration is what makes our targeted approach so efficient. Similarly, the near-perfect predictive power of the dual-signal model (reducing mean absolute error by half compared to confidence-only) provides a reliable foundation for implementing quality control mechanisms in qualitative research at scale.
Methodological Significance and Innovation
The dual-signal calibration approach represents a novel paradigm for model calibration in qualitative research. While previous work has primarily relied on single metacognitive signals such as confidence, our integration of model diversity provides a more robust framework for reliability assessment. This builds upon theoretical foundations in ensemble learning, where diversity has long served as an error bound, but extends this concept to qualitative coding in a novel application. The optimized risk score formula balances these signals based on empirical evidence, creating a transparent and adjustable mechanism for risk assessment. Our emphasis on diversity resonates with deep‐ensemble strategies … 21, 29, 25 and complements post-hoc reliability metrics such as the trust score proposed by Jiang et al.23, thereby bridging machine-learning calibration research with qualitative methodology.
Practical Applications and Implementation Guidelines
Implementation is straightforward: gather confidence ratings from 4–8 models, compute diversity (normalised entropy) and the combined risk score . Items with are auto-accepted, receive a cursory check, and undergo full review. For a 1,000-segment study this triage removes roughly 3,250 manual checks, saving two working days of expert time. Thresholds can be tightened for high-stakes contexts or relaxed when resources are limited.
The high predictive capacity of our dual-signal model () allows researchers to reliably forecast inter-model agreement. In practical terms, qualitative researchers can quantitatively estimate coding reliability without extensive human validation or ground truth labels—a significant advancement for resource-constrained projects. This 50% MAE reduction roughly halves misclassification and can save many hours of manual correction on large projects.
The practical implications of the error concentration pattern (87% of errors in 50% of data) are profound for qualitative research practice. First, researchers can confidently auto-accept over one-third of coding decisions without meaningful quality loss, allowing reallocation of up to 35% of coding budget toward more interpretive tasks. Second, a positive feedback loop emerges: as experts focus on high-risk segments, their attention deepens and overall coding quality improves.
Specifically, implementation unfolds in three stages. First, researchers collect coding decisions together with five-point confidence ratings from a panel of four to eight models. Second, they compute the normalised entropy–based diversity index for every coding point and combine it with mean confidence to obtain the joint risk score . Third, the score is used to automatically accept low-risk items (), flag medium-risk items () for cursory inspection, and route high-risk items () to full expert review. This compact procedure preserves the logic of the workflow while eliminating the need for bullet-point instructions.
This structured approach maximizes the return on expert time investment while maintaining rigorous quality standards, a critical balance in large-scale qualitative research projects.
Implementation Checklist
To facilitate practical adoption, we distil the above workflow into the following six-step checklist:
1.
Assemble panel — Choose 4–8 diverse LLMs; fix generation settings (e.g., temperature=0.2).
2.
Collect signals — For every coding point, record each model’s categorical vote (single label) and a normalised confidence score in (e.g., scale 1–5 ratings or use softmax probability).
3.
Compute metrics — Derive mean confidence and diversity (normalised entropy). See Supplementary Algorithm 1.
4.
Risk scoring — Calculate .
5.
Route decisions — Apply thresholds: → auto-accept; → light review; → full review.
6.
Audit & archive — Log prompts and outputs; audit per tier to monitor drift.
Supplementary Table 7 summarises recommended cut-offs and expected review load for three common risk budgets, while a printable six–step reference is provided in Supplementary Table 9.
Threshold tips: In high–risk domains (e.g., medicine, law) tighten the green-zone cut-off to , reducing residual error to while increasing human review by 10 pp.
Limitations and Boundary Conditions
Several limitations warrant consideration. First, diversity measurement has reduced resolution in smaller panels (e.g., four‐model panels), limiting the granularity of risk assessment14 (see Supplementary Table 6 and Extended Data Fig. 6). Second, the dual–signal regression in Fig. 1 is fitted on ten category–level observations, providing an illustrative but low-powered test (). Bootstrap validation (Supplementary Table 11) confirms the robustness of the diversity coefficient, yet the wide intervals for the intercept and confidence term reflect this sample constraint. Nevertheless, the six-dataset cross-domain validation (Results; Extended Data Fig. 7) reproduces the same directional pattern, substantially mitigating concerns about limited statistical power. An item-level mixed-effects re-analysis (see Supplementary Table 12) yields similar slopes but narrower intervals, supporting our qualitative conclusions. Third, for the binary Finance–News benchmark the baseline agreement is already near ceiling (83 % unanimous labels), which mechanically caps explainable variance such that the dual-signal model raises only marginally (0.1460.154). We interpret this as a ceiling effect rather than a failure of the method. Fourth, our approach is validated on relatively straightforward coding tasks with conceptually accessible categories. More complex, nuanced coding scenarios—such as highly abstract theoretical constructs, deeply contextual interpretations, or tasks requiring extensive domain expertise—may require different calibration approaches and warrant separate investigation. Fifth, conceptual clarity significantly impacts method effectiveness, with abstract categories requiring more human intervention regardless of confidence levels. Finally, language and cultural specificity may present challenges for cross-lingual applications that require further investigation.
Our primary corpus comprises 71 interview segments from four participants and therefore illustrates methodological feasibility rather than population inference. Replication on six public benchmark datasets mitigates this limitation but does not replace in-domain validation. Ongoing data collection will broaden both the size and diversity of the interview corpus, enabling more robust tests of external validity in future work.
Integration of Theory and Practice
The dual-signal mechanism connects to broader theories in cognitive science regarding metacognition and judgment calibration. It reframes the role distribution in human-AI collaborative coding from exhaustive human review to targeted expert intervention where most needed. This represents a new perspective on balancing scalability and reliability in qualitative research, complementing other assistive technologies such as topic modeling and automatic summarization while maintaining the interpretive depth characteristic of qualitative inquiry.
Future Research Directions
Future work should extend this approach to additional domains and task types, including multi-label classification, sequence labeling, cross-cultural qualitative analysis, and more complex interpretive frameworks.30 A particularly important direction involves investigating how confidence-diversity calibration performs on highly complex coding tasks that require deep contextual understanding, abstract theoretical reasoning, or extensive domain expertise. Such tasks may necessitate modified calibration approaches, additional metacognitive signals, or hybrid human-AI workflows that differ from our current three-tier system. Refinements to the risk scoring formula could incorporate additional metacognitive signals or adjust weights based on domain-specific requirements. Dynamic threshold systems that automatically adjust based on real-time feedback represent another promising direction. Additionally, developing specialized expert-AI interaction interfaces could further optimize human-machine collaborative efficiency.
Broader Implications and Ethical Considerations
Our approach lowers the cost of large-scale qualitative analysis, but also amplifies the ethical stakes of automated interpretation. Transparent prompts, versioned model checkpoints and public error audits are therefore essential safeguards 23, 31. Moreover, governance frameworks must recognise that calibration can mitigate but not eliminate demographic bias inherited from training data; independent human review remains indispensable for sensitive domains such as health and justice 24.
Conclusion
The dual-signal calibration method provides a reliable framework for AI-assisted coding in qualitative research, while the three-tier workflow significantly improves efficiency without compromising quality. The cross-domain applicability offers researchers across disciplines a practical tool for enhancing their methodological toolkit. Human-AI collaborative coding represents an important methodological advancement that respects the interpretive nature of qualitative research while addressing its scalability challenges.
Methods
Data sources and ethical considerations
The primary corpus comprised 71 text segments drawn from semi-structured interviews with four adult participants selected from a larger longitudinal study on academic identity in digital media. The study protocol was approved by the Institutional Review Board of South China University of Technology (IRB# 2024–Q–017), and all interviewees provided written informed consent allowing secondary analysis. To assess the external validity of our approach we additionally evaluated six open, task-diverse benchmark datasets—Finance-News600, PubMed-RCT600, ECHR-A, Moral-Stories, WMT20 QE and XNLI-zh 32, 33, 34, 35, 36. Dataset licences and digital object identifiers are listed in Supplementary Table 1.
Computational models and environment
We queried eight large language models (LLMs) representative of three vendor families: Claude 3.5 & 3.7 Sonnet (Anthropic, 2024–05–28 release), GPT–4o and GPT–o3-Mini (OpenAI, 2024–05–13), Gemini 2.5 Flash (Google, 2024–06–01) in single-response and chain-of-thought variants, and DeepSeek V3 & Reasoner (DeepSeek, 2024–05–20). All API calls were issued via HTTPS with temperature (single response) or (chain-of-thought), top_p and max_tokens . Inference ran on an Ubuntu 22.04 server equipped with four NVIDIA A100 80 GB GPUs, Python 3.10.13, PyTorch 2.1.0 and HuggingFace transformers 4.40. Random seeds were fixed to 42 and every request–response pair was logged to enable exact reproduction.
Dual-signal calibration framework
For each model–segment–category triple we recorded the binary coding decision, a self-reported confidence score and, when permitted, the full chain-of-thought trace. Reliability was characterised by two orthogonal signals: (i) the mean confidence across the eight models; (ii) decision diversity , computed as the normalised Shannon entropy of the vote distribution. These quantities were combined into a risk score
where lower values denote higher expected reliability. Coding points were automatically routed to one of three review tiers—auto-accept, light review or full review—using empirically determined cut-offs (Extended Data Fig. 5).
To determine the relative weight of the confidence term we conducted an exhaustive grid search over in the generic form . Ten-fold cross-validation on all segment–code units minimised the mean absolute error at (pure diversity, ). A secondary plateau was observed for (, ), which delivered markedly stronger performance on six out-of-domain datasets (median ). Balancing interpretability, internal fit and cross-domain robustness, we therefore lock for the remainder of the study; the full –error curve is provided in Extended Data Fig. 7 (Supplementary Materials). Per-category cross-validation metrics are summarised in Supplementary Table 8.
Human validation procedure
A stratified random sample of coding points from each review tier (total ) was independently annotated by three trained qualitative researchers who were blinded to model identity (Cohen’s ). Items lacking a two-coder majority proceeded to expert adjudication by two senior scholars in qualitative methodology. The adjudicated label served as ground truth for evaluating accuracy, precision, recall and scores. The green-zone residual-error rate reported in the Results section is derived from this audit: it equals the fraction of audited auto-accepted points whose AI label disagreed with expert consensus, extrapolated to the full green zone.
Statistical analysis
All analyses were conducted in Python 3.10 using pandas 2.2, statsmodels 0.14 and SciPy 1.12, and correlation coefficients are reported as Pearson’s 37. Ninety-five-per-cent confidence intervals were obtained via -fold non-parametric bootstrap over interview segments. Linear and multiple regressions were fit with ordinary least squares; predictive performance was assessed with nested ten-fold cross-validation using mean absolute error (MAE) and . Correlation coefficients are reported as Pearson’s . Hypothesis tests were two-sided with , and we controlled the false discovery rate using Benjamini–Hochberg correction. Power analysis conducted in G*Power 3.1 indicated that the human-validation sample affords power to detect a percentage-point difference in accuracy between tiers.
Coding framework
All analyses rely on the ten-theme coding scheme originally developed for the “Academic Identity and Knowledge Strategies” project. Full operational definitions and exemplar quotes are provided in Supplementary Table 2; the plain-language label set is archived with the code repository. During inference each LLM received the category name, its two-sentence definition, and an abridged decision rule. Prompts and exact JSON payloads are included in the public release (see Code availability).
Code availability
The complete inference and analysis pipeline is publicly available at Harvard Dataverse (DOI 10.7910/DVN/G1AYGK). The reproducibility package includes all analysis scripts, plotting code, and a one-click reproduction script (run_all.sh) that regenerates all main figures and results. All notebooks and scripts reproduce the figures from raw model outputs.
Data availability
The replication data package is publicly available at Harvard Dataverse under the title “Replication Data for ‘Confidence–Diversity Calibration of AI Judgement Enables Reliable Qualitative Coding”’ (DOI 10.7910/DVN/G1AYGK). The package includes processed data for all eight main models, extracted results in CSV format, and all figures and tables. An anonymised subset of interview segments sufficient to reproduce the main findings is included in the package. The complete, identifiable corpus can be provided by the corresponding author under IRB-approved data-use agreements. External validation datasets are publicly accessible via the links in Supplementary Table 1. The data package is released under Creative Commons CC0 1.0 Universal Public Domain Dedication.
Baseline human–only coding
Two trained research assistants (Coder 1 and Coder 2) independently labelled all 71 interview segments against the ten-theme scheme, yielding segment–code units and therefore binary decisions. These labels constitute the conventional two-coder baseline reported in Table 1. Cohen’s and the percentage error were computed with the scikit--learn implementation of cohen_kappa_score, where error is defined as the proportion of non-identical labels between the two coders. The third human coder was reserved for the later expert adjudication step and is not included in this baseline analysis.
Reporting summary. Further information on research design is available in the Nature Portfolio Reporting Summary linked to this article.
Supplementary information The online version contains supplementary material available at https://doi.org/10.7910/DVN/G1AYGK.
Acknowledgements
The authors gratefully acknowledge the financial support of the Fundamental Research Funds for the Central Universities project “Innovation in Journalism and Communication in the Age of Intelligence” (grant XHJH202504). We thank Ms Wenqing Tang (School of Journalism and Communication, South China University of Technology) for collecting the interview data that form the primary corpus analysed in this study, and we are indebted to all interview participants for their time and insights. Computational experiments and large–language–model inference were performed on the High–Performance Computing Platform of the Guangdong–Hong–Kong–Macao Greater Bay Area Research Institute of International Communication, South China University of Technology. All procedures were approved by the Institutional Review Board of South China University of Technology (approval SCUTXWWX230057).
Author Contributions
Z.Z. conceived the study, secured funding, designed the methodology, developed the software, performed formal analysis and visualisation, and wrote the original draft. Y.L. collected and curated the data, reviewed and edited the manuscript, supervised the project and managed project administration. Both authors read and approved the final manuscript.
Declaration of Interests
The authors declare no competing interests.
Figure & Table Legends
Table 1: Reliability of human versus AI coding. Comparison of Cohen’s , error rate and number of decisions across the conventional two-coder baseline, the eight-model AI majority vote, and the full three-tier workflow integrating expert review.
Figure 1: Dual-signal mechanism for calibrating AI qualitative coding. A Mean self-confidence versus full agreement (linear regression , ). B Residuals plotted against model diversity. C Two-dimensional confidence–diversity plane (multiple regression, ); colour scale denotes observed agreement. Equations and values are annotated within the panels for quick reference.
Figure 2: Three-tier workflow optimizes human-AI collaboration. A Confidence–diversity plane showing the partitioning into three actionable zones: green (auto-accept), amber (light audit), and red (full review). B Distribution of segments and errors across the three workflow tiers, showing error concentration in the full review tier. C Coding point-level risk distribution based on the quantitative risk score (Eq. 1), with horizontal line indicating the proportion of points selected for expert audit. D Efficiency-accuracy trade-off curve showing pass rates and error rates at different confidence thresholds, illustrating how higher thresholds reduce errors but require more manual review.
Figure 3: Prompting style and model family comparison. Cohen’s and confidence for single vs. CoT variants across four model families.
Figure 4: External validation across six public datasets. A Baseline and enhanced Cohen’s . B Baseline versus enhanced .
Figure 5: Three-Tier Workflow Decision Tree. Risk score derived from mean confidence () and diversity () routes each coding point to auto-accept (green), light review (amber) or full expert review (red), maximising efficiency while maintaining reliability.